# FancyVideo: Towards Dynamic and Consistent Video Generation via Cross-frame Textual Guidance

### v2

@@ Abstract
Synthesizing motion-rich and temporally consistent videos remains a challenge in artificial intelligence, especially when dealing with extended durations. Existing text-to-video (T2V) models commonly employ spatial cross-attention for text control, equivalently guiding different frame generations without frame-specific textual guidance. Thus, the model’s capacity to comprehend the temporal logic conveyed in prompts and generate videos with coherent motion is restricted. To tackle this limitation, we introduce FancyVideo, an innovative video generator that improves the existing text-control mechanism with the well-designed Cross-frame Textual Guidance Module (CTGM). Specifically, CTGM incorporates the Temporal Information Injector (TII), Temporal Affinity Refiner (TAR), and Temporal Feature Booster (TFB) at the beginning, middle, and end of cross-attention, respectively, to achieve frame-specific textual guidance. Firstly, TII injects frame-specific information from latent features into text conditions, thereby obtaining cross-frame textual conditions. Then, TAR refines the correlation matrix between cross-frame textual conditions and latent features along the time dimension. Lastly, TFB boosts the temporal consistency of latent features. Extensive experiments comprising both quantitative and qualitative evaluations demonstrate the effectiveness of FancyVideo. Our approach achieves state-of-the-art T2V generation results on the EvalCrafter benchmark and facilitates the synthesis of dynamic and consistent videos. Our video demo, code and model are available at https://360cvgroup.github.io/FancyVideo/.
@@
11footnotetext: These authors contributed equally.22footnotetext: Corresponding Authors.
Introduction
With the advancement of the diffusion model, the text-to-image (T2I) generative models (Blattmann et al. 2023b; Ge et al. 2023; Ho et al. 2022; Luo et al. 2023) can produce high-resolution and photo-realistic images by complex text prompts, resulting in various applications. Currently, many studies (Yu et al. 2023; Guo et al. 2023a; Wang et al. 2024; Zhou et al. 2022) explore the text-to-video (T2V) generative model due to the great success of T2I models. However, building a powerful T2V model remains challenging as it requires maintaining temporal consistency while generating coherent motions simultaneously. Moreover, due to limited memory, most diffusion-based T2V models (Yu et al. 2023; Guo et al. 2023a; Zhang et al. 2024; Guo et al. 2023b; Chen et al. 2023; Menapace et al. 2024; Wang et al. 2023b) can only produce fewer than 16 frames of video per sampling without extra assistance (i.e., super-resolution).
The existing T2V models (Zhang et al. 2024; Guo et al. 2023b; Chen et al. 2023; Menapace et al. 2024; Wang et al. 2023b) typically employ spatial cross-attention between text conditions and latent features for achieving text control generation. However, as shown in Fig. 2(I), this manner shares the same text condition across different frames, thus lacking the specific textual guidance tailored to each frame. Consequently, these T2V models struggle to comprehend the temporal logic of text prompts and produce videos with coherent motion. Taking AnimateDiff (Guo et al. 2023b) as an example, in Fig. 1, we exhibit its generated video and visualize the [verb]-focused region (which is closely associated with the video motion) based on the attention map from the cross-attention module. Ideally, these regions should transition smoothly over time and align with the semantics of motion instructions. However, as observed in the upper right of the figure, the [verb]-focused region remains nearly identical across different frames due to the consistent textual guidance between frames. Meanwhile, the video exhibits poor motion in the upper left of the figure. Furthermore, we perform a similar visual analysis for the longer video (e.g., 64 frames) generation and find that this problem is more prominent, as illustrated in the lower part of Fig. 1. Therefore, we believe this approach hampers the advancement of video dynamics and consistency and is sub-optimal for video generation tasks based on text prompts.
@@ Introduction
To this end, we present a novel T2V model named FancyVideo, capable of comprehending complex spatial-temporal relationships within text prompts. By employing a cross-frame textual guidance strategy, FancyVideo can generate more dynamic and plausible videos in a sampling process. Specifically, to boost the model’s capacity for understanding spatial-temporal information in text prompts, we optimize the spatial cross-attention through the proposed Cross-frame Textual Guidance Module (CTGM), comprising a Temporal Information Injector (TII), Temporal Affinity Refiner (TAR), and Temporal Feature Booster (TFB). As illustrated in Fig. 2(II), TII injects temporal information from latent features into text conditions, building cross-frame textual conditions. Then, TAR refines the affinity between frame-specific text embedding and video along time dimension, adjusting the temporal logic of textual guidance. Lastly, TFB further boosts the temporal consistency of latent features. Through the cooperative interaction between TII, TAR, and TFB, FancyVideo fully captures the motion logic embedded within images and text. Consequently, its motion token-focused area shifts logically with frames, as illustrated in the lower right part of Fig. 1. This characteristic enables FancyVideo to produce dynamic videos, as displayed in the lower left part of the figure. Experiments demonstrate that FancyVideo successfully generates dynamic and consistent videos, achieving the SOTA results on the EvalCrafter (Liu et al. 2023) benchmark and the competitive performance on UCF-101 (Soomro, Zamir, and Shah 2012) and MSR-VTT (Xu et al. 2016).
@@
Research in this field has employed various generative models, including GANs (Zhu et al. 2017; Wang et al. 2020; Munoz et al. 2021; Gur, Benaim, and Wolf 2020), auto-regressive models (Wang et al. 2019; Ramesh et al. 2021; Yan et al. 2021; Razavi, Van den Oord, and Vinyals 2019), and implicit neural representations (De Luigi et al. 2023). Recently, diffusion models (Zhang, Rao, and Agrawala 2023; Rombach et al. 2022; Zhang et al. 2023a; Ruiz et al. 2023) have advanced text-to-image generation quality, outperforming previous techniques. Stable Diffusion (Rombach et al. 2022) adapts the diffusion process to VAE’s (Kingma and Welling 2013) latent space, reducing computation and memory requirements. Inspired by T2I success, T2V studies (Ren et al. 2024; Jiang et al. 2023; Xing et al. 2024; Wu et al. 2023a) add temporal layers to pre-trained T2I models to capture temporal information while reducing video generation costs. However, these methods often use a single text condition for spatial cross-attention, leading to inconsistent temporal context across frames and affecting motion coherence. Our work introduces a cross-frame textual guidance mechanism to tailor temporal context for each frame, generating more realistic and temporally consistent videos.
@@
The gap between text prompts and videos has steered research towards using images for clearer video generation. SVD (Blattmann et al. 2023a) uses images as noisy latent variables for frame generation, while MoonShot (Zhang et al. 2024) enhances semantic consistency through the CLIP encoder. These I2V methods require specific input images. Hierarchical approaches (Zeng et al. 2023; Chen et al. 2023) use images as key frames, potentially creating lengthy videos with fewer image constraints. These methods, essentially T2V, also support I2V. FancyVideo, a hierarchical method with cross-frame textual guidance, generates more frames per iteration, cutting down inference time.
Method
Preliminaries
@@
The latent diffusion models (LDMs) (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) are a class of efficient diffusion models that convert the denoising process into the compressed latent space instead of the pixel space. Specifically, LDMs employ VAE’s (Kingma and Welling 2013) encoder to compress the image into latent code and learn the data distribution by performing a forward and inverse diffusion process on the latent code. It assumes a forward process that gradually introduces the gaussian noise ($\epsilon\sim\mathcal{N}(0,I)$) into latent code ($\mathbf{z}$), getting:
@@
where $\mathbf{c}$ represents the condition and target $\boldsymbol{y}$ can be noise $\epsilon$, denoising input $\mathbf{z}$ or $\boldsymbol{v}$-prediction ($\boldsymbol{v}=\sqrt{\bar{\alpha_{t}}}\epsilon-\sqrt{1-\bar{\alpha_{t}}}\mathbf{z}$) in (Salimans and Ho 2022). In this paper, we adopt the $\boldsymbol{v}$-prediction as the supervision.
@@
Previous studies proposed zero terminal SNR (Lin et al. 2024) to handle the signal-to-noise ratio (SNR) difference between the testing and training phase, which hinders the generation quality. At training, due to the residual signal left by the noise scheduler, the SNR is still not zero at the terminal timestep $T$. However, the sampler lacks realistic data when sampling from random gaussian noise during the test, resulting in a zero SNR. This train-test discrepancy is unreasonable and an obstacle to generating high-quality videos. Therefore, following the (Lin et al. 2024; Gupta et al. 2023; Girdhar et al. 2023), we scale up the noise schedule and set $\bar{\alpha_{T}}=0$ to fix this problem.
Model Architecture
Figure 3: The overall architecture of our method. FancyVideo is a T+I2V model that concatenates noise latent, mask indicator, and image indicator as input. We insert our Cross-frame Textual Guidance Module (CTGM) into each spatial block. CTGM consists of three components: Temporal Information Injector, Temporal Affinity Refiner, and Temporal Feature Booster. These components are inserted at the beginning, middle, and end of cross-attention, respectively.
@@
Fig. 3 illustrates the overall architecture of FancyVideo. The model is structured as a pseudo-3D UNet, which integrates frozen spatial blocks, sourced from a text-to-image model, along with Cross-frame Textual Guidance Modules (CTGM) and temporal attention blocks. The model takes three features as input: noisy latent $\mathcal{Z}_{n}\in\mathbb{R}^{f\times h\times w\times c}$, where $h$ and $w$ indicate the height and width of the latent, $f$ signifies the number of frames, and $c$ denotes the channels of the latent; mask indicator $\mathcal{M}\in\mathbb{R}^{f\times h\times w\times 1}$, with elements set to 1 for the first frame and 0 for all other frames; image indicator $\mathcal{I}\in\mathbb{R}^{f\times h\times w\times c}$, with initial image as the first frame and 0 for all other frames. The denoising input $\mathcal{Z}$ is formed by concatenating $\mathcal{Z}_{n}$, $\mathcal{M}$ and $\mathcal{I}$ along the channel dimension, represented as $\mathcal{Z}=[\mathcal{Z}_{n};\mathcal{M};\mathcal{I}]\in\mathbb{R}^{f\times h\times w\times(2c+1)}$. Within each spatial block, CTGM is employed to capture the intricate dynamics described in the text prompts. Afterward, we apply temporal attention blocks to enhance the temporal relationships across various patches.
Cross-frame Textual Guidance Module
@@
CTGM advances the existing text control method through three sub-modules: Temporal Information Injector (TII), Temporal Affinity Refiner (TAR), and Temporal Feature Booster (TFB) as depicted in Fig. Method(III). Before engaging in cross-attention, TII initially extracts temporal latent feature $\mathcal{Z}_{t}$ and then incorporates temporal information into text embedding $\mathcal{T}_{rep}$ based on $\mathcal{Z}_{t}$, obtaining cross-frame textual condition $\mathcal{T}_{z}$. Subsequently, TAR refines the affinity between $\mathcal{Z}_{t}$ and $\mathcal{T}_{z}$ along the time axis, enhancing the temporal coherence of textual guidance. Lastly, TFB further enhances the temporal continuity of the feature. The computation process of the CTGM can be formalized as:
@@
$\mathcal{Z}^{\prime}_{ref}=\mathrm{TFB}(\mathrm{Softmax}($ $\frac{\mathrm{TAR}(W_{q}\mathcal{Z}_{t},W_{k}\mathcal{T}_{z})}{\sqrt{d_{k}}}W_{v}(\mathcal{T}_{z})),$ (4)
where $W_{q}$, $W_{k}$, and $W_{v}$ represent the linear layers for query, key, and value in original cross-attention, respectively. The hyper-parameter $d_{k}$ is acquired from the query dimensions. $\mathrm{TII}(\cdot,\cdot)$, $\mathrm{TAR}(\cdot)$, $\mathrm{TFB}(\cdot)$ denotes the TII, TAR and TFB. In the end, we get refined noisy latent feature ${Z}^{\prime}_{ref}$. A detailed description of these three modules is provided as follows.
@@
In previous work (Guo et al. 2023b; Girdhar et al. 2023), the text embedding $\mathcal{T}_{rep}$ is repeated equally $f$ times, resulting in $\mathcal{T}_{rep}\in\mathbb{R}^{f\times n\times c}$, $n$ denoting the length of the embedding vector. We inject temporal information into the embedding before performing spatial cross-attention, thereby enabling distinct focal points on the text within different frames. In Temporal Information Injector (TII), we initially reshape the noisy latent $\mathcal{Z}$ from $\mathbb{R}^{f\times h\times w\times c}$ to $\mathbb{R}^{(hw)\times f\times c}$ and apply temporal self-attention to acquire $\mathcal{Z}_{t}$. Then, we conduct spatial cross-attention, using the repeated text embedding $\mathcal{T}_{rep}$ as queries and the noisy latent $\mathcal{Z}_{t}\in\mathbb{R}^{f\times(hw)\times c}$ as both keys and values, resulting in the text embedding $\mathcal{T}_{z}$ with frame-specific temporal information. The formalization of the TII module can be expressed as follows:
@@
Temporal Feature Booster.
Figure 4: Qualititive analysis. We compare the video generation results from AnimateDiff (Guo et al. 2023b), DynamiCrafter (Xing et al. 2023), Pika (PikaLabs 2024), Gen-2 (Runway 2024), and our FancyVideo.
Dimensions Metrics Pika(2024) Gen2(2024) Show-1(2023b) Model- Scope(2023a) Dynami- Crafter(2023) Animate- Diff(2023b) FancyVideo
Video Quality VQAA($\uparrow$) 59.09 59.44 23.19 40.06 74.56 65.94 85.78
VQAT($\uparrow$) 64.96 76.51 44.24 32.93 59.48 52.02 74.56
IS($\uparrow$) 14.81 14.53 17.65 17.64 18.37 16.54 17.38

### v4

@@ Abstract
Synthesizing motion-rich and temporally consistent videos remains a challenge in artificial intelligence, especially when dealing with extended durations. Existing text-to-video (T2V) models commonly employ spatial cross-attention for text control, equivalently guiding different frame generations without frame-specific textual guidance. Thus, the model’s capacity to comprehend the temporal logic conveyed in prompts and generate videos with coherent motion is restricted. To tackle this limitation, we introduce FancyVideo, an innovative video generator that improves the existing text-control mechanism with the well-designed Cross-frame Textual Guidance Module (CTGM). Specifically, CTGM incorporates the Temporal Information Injector (TII) and Temporal Affinity Refiner (TAR) at the beginning and end of cross-attention, respectively, to achieve frame-specific textual guidance. Firstly, TII injects frame-specific information from latent features into text conditions, thereby obtaining cross-frame textual conditions. Then, TAR refines the correlation matrix between cross-frame textual conditions and latent features along the time dimension. Extensive experiments comprising both quantitative and qualitative evaluations demonstrate the effectiveness of FancyVideo. Our approach achieves state-of-the-art T2V generation results on the EvalCrafter benchmark and facilitates the synthesis of dynamic and consistent videos. Note that the T2V process of FancyVideo essentially involves a text-to-image step followed by T+I2V. This means it also supports the generation of videos from user images, i.e., the image-to-video (I2V) task. A significant number of experiments have shown that its performance is also outstanding.
footnotetext: Equal contribution.footnotetext: Project leader.footnotetext: Corresponding author.
1 Introduction
With the advancement of the diffusion model, the text-to-image (T2I) generative models Blattmann et al. (2023b); Ho et al. (2022); Luo et al. (2023); Ma et al. (2024); Liu et al. (2025); Cao et al. (2025); Ling et al. (2025); Bi et al. (2024); Ma et al. (2025) can produce high-resolution and photo-realistic images by complex text prompts, resulting in various applications. Currently, many studies Wang et al. (2024); Guo et al. (2023a) explore the text-to-video (T2V) generative model due to the great success of T2I models. However, building a powerful T2V model remains challenging as it requires maintaining temporal consistency while generating coherent motions simultaneously. Moreover, due to limited memory, most diffusion-based T2V models Wang et al. (2024); Guo et al. (2023a); Zhang et al. (2024a); Guo et al. (2023b); Chen et al. (2023); Menapace et al. (2024) can only produce fewer than 16 frames of video per sampling without extra assistance (i.e., super-resolution).
The existing T2V models Zhang et al. (2024a); Guo et al. (2023b); Chen et al. (2023); Menapace et al. (2024); Bi et al. (2025); Wang et al. (2025a); Shao et al. (2025) typically employ spatial cross-attention between text conditions and latent features for achieving text control generation. However, as shown in Fig. 2(I), this manner shares the same text condition across different frames, thus lacking the specific textual guidance tailored to each frame. Consequently, these T2V models struggle to comprehend the temporal logic of text prompts and produce videos with coherent motion. Taking AnimateDiff Guo et al. (2023b) as an example, in Fig. 1, we exhibit its generated video and visualize the [verb]-focused region (which is closely associated with the video motion) based on the attention map from the cross-attention module. Ideally, these regions should transition smoothly over time and align with the semantics of motion instructions. However, as observed in the upper right of the figure, the [verb]-focused region remains nearly identical across different frames due to the consistent textual guidance between frames. Meanwhile, the video exhibits poor motion in the upper left of the figure.
@@ 1 Introduction
Furthermore, we perform a similar visual analysis for the longer video (e.g., 64 frames) generation and find that this problem is more prominent, as illustrated in the lower part of Fig. 1. Therefore, we believe this approach hampers the advancement of video dynamics and consistency and is sub-optimal for video generation tasks based on text prompts.
To this end, we present a novel T2V model named FancyVideo, capable of comprehending complex spatial-temporal relationships within text prompts. By employing a cross-frame textual guidance strategy, FancyVideo can generate more dynamic and plausible videos in a sampling process. Specifically, to boost the model’s capacity for understanding spatial-temporal information in text prompts, we optimize the spatial cross-attention through the proposed Cross-frame Textual Guidance Module (CTGM), comprising a Temporal Information Injector (TII) and Temporal Affinity Refiner (TAR). As illustrated in Fig. 2(II), TII injects temporal information from latent features into text conditions, building cross-frame textual conditions. Then, TAR refines the affinity between frame-specific text embedding and video along time dimension, adjusting the temporal logic of textual guidance. Through the cooperative interaction between TII and TAR, FancyVideo fully captures the motion logic embedded within images and text. Consequently, its motion token-focused area shifts logically with frames, as illustrated in the lower right part of Fig. 1. This characteristic enables FancyVideo to produce dynamic videos, as displayed in the lower left part of the figure. Experiments demonstrate that FancyVideo successfully generates dynamic and consistent videos, achieving the SOTA results on the EvalCrafter Liu et al. (2023) benchmark and the competitive performance on UCF-101 Soomro et al. (2012) and MSR-VTT Xu et al. (2016). Additionally, FancyVideo supports generating videos from user-input images, i.e., the image-to-video task. We have also conducted extensive experiments to demonstrate the superiority of our method.
@@ 2 Related Work
Generative models like GANs Wang et al. (2020); Munoz et al. (2021); Gur et al. (2020), auto-regressive models Wang et al. (2019); Yan et al. (2021), and implicit neural representations De Luigi et al. (2023) have been explored for video generation. Recently, diffusion models Rombach et al. (2022); Zhang et al. (2024b, c) have advanced text-to-image quality. Stable Diffusion Rombach et al. (2022) uses a VAE Kingma and Welling (2013) latent space to reduce cost Jiang et al. (2023). T2V models Wu et al. (2023a) add temporal layers to T2I models but often lack frame-to-frame consistency. We propose cross-frame textual guidance to improve temporal coherence.
@@
To bridge the gap between text and video, recent work leverages images for clearer video generation. SVD Blattmann et al. (2023a) treats images as noisy latent inputs, while MoonShot Zhang et al. (2024a) improves semantic consistency using a CLIP encoder. Though effective, these I2V methods rely on input images. Hierarchical approaches Zeng et al. (2023); Chen et al. (2023) use images as keyframes to extend video length with fewer constraints. These methods, though I2V-capable, are essentially T2V. FancyVideo adopts a hierarchical design with cross-frame textual guidance, enabling more frames per iteration and faster inference.
3 Method
Figure 3: The overall architecture of our method. FancyVideo is a T+I2V model that concatenates noise latent, mask indicator, and image indicator as input. We insert our Cross-frame Textual Guidance Module (CTGM) into each spatial block. CTGM consists of three components: Temporal Information Injector, Temporal Affinity Refiner, and Temporal Feature Booster (see supplementary materials). These components are inserted at the beginning, middle, and end of cross-attention, respectively.
3.1 Preliminaries
@@
LDMs Sohl-Dickstein et al. (2015); Ho et al. (2020); Zhang et al. (2025); He et al. (2025); Lu et al. (2025) enhance efficiency by running diffusion in the VAE-compressed latent space Kingma and Welling (2013) instead of pixel space. The forward process adds Gaussian noise ($\epsilon\sim\mathcal{N}(0,I)$) to the latent code $\mathbf{z}$, yielding:
@@
where $\mathbf{c}$ represents the condition and target $\boldsymbol{y}$ can be noise $\epsilon$, denoising input $\mathbf{z}$ or $\boldsymbol{v}$-prediction ($\boldsymbol{v}=\sqrt{\bar{\alpha_{t}}}\epsilon-\sqrt{1-\bar{\alpha_{t}}}\mathbf{z}$) in Salimans and Ho (2022). In this paper, we adopt the $\boldsymbol{v}$-prediction as the supervision.
@@
Previous studies proposed zero terminal SNR Lin et al. (2024) to handle the signal-to-noise ratio (SNR) difference between the testing and training phase, which hinders the generation quality. At training, due to the residual signal left by the noise scheduler, the SNR is still not zero at the terminal timestep $T$. However, the sampler lacks realistic data when sampling from random gaussian noise during the test, resulting in a zero SNR. This train-test discrepancy is unreasonable and an obstacle to generating high-quality videos. Therefore, following the Lin et al. (2024); Girdhar et al. (2023), we scale up the noise schedule and set $\bar{\alpha_{T}}=0$ to fix this problem.
3.2 Model Architecture
@@
Fig. 3 illustrates the overall architecture of FancyVideo. The model is structured as a pseudo-3D UNet, which integrates frozen spatial blocks, sourced from a text-to-image model, along with Cross-frame Textual Guidance Modules (CTGM) and temporal attention blocks. The model takes three features as input: noisy latent $\mathcal{Z}_{n}\in\mathbb{R}^{f\times h\times w\times c}$, where $h$ and $w$ indicate the height and width of the latent, $f$ signifies the number of frames, and $c$ denotes the channels of the latent; mask indicator $\mathcal{M}\in\mathbb{R}^{f\times h\times w\times 1}$, with elements set to 1 for the first frame and 0 for all other frames; image indicator $\mathcal{I}\in\mathbb{R}^{f\times h\times w\times c}$, with initial image as the first frame and 0 for all other frames. The denoising input $\mathcal{Z}$ is formed by concatenating $\mathcal{Z}_{n}$, $\mathcal{M}$ and $\mathcal{I}$ along the channel dimension, represented as $\mathcal{Z}=[\mathcal{Z}_{n};\mathcal{M};\mathcal{I}]\in\mathbb{R}^{f\times h\times w\times(2c+1)}$. Within each spatial block, we first incorporate prior knowledge of the motion score as embeddings. In each subsequent cross-attention layer, CTGM is employed to capture the intricate dynamics described in the text prompts. Afterward, we apply temporal attention blocks to enhance the temporal relationships across various patches.
3.2.1 Motion Embedding
To achieve more controllable video generation in terms of motion amplitude, we introduce motion score information calculated by the RAFT Teed and Deng (2020) alongside the timestep information. Specifically, we calculate a motion score for the training samples in the dataset within a range of 0.1 to 10. The score are then encoded into motion features through a motion embedding layer. By controlling the motion score, we can generate videos with stronger motion. However, simply adjusting the score may lead to unrealistic motion. We use CTGM to prevent these issues.
3.2.2 Cross-frame Textual Guidance Module
@@
CTGM advances the existing text control method through two sub-modules: Temporal Information Injector (TII) and Temporal Affinity Refiner (TAR) as depicted in Fig. 3(III). Before engaging in cross-attention, TII initially extracts temporal latent feature $\mathcal{Z}_{t}$ and then incorporates temporal information into text embedding $\mathcal{T}_{rep}$ based on $\mathcal{Z}_{t}$, obtaining cross-frame textual condition $\mathcal{T}_{z}$. Subsequently, TAR refines the affinity between $\mathcal{Z}_{t}$ and $\mathcal{T}_{z}$ along the time axis, enhancing the temporal coherence of textual guidance. The computation process of the CTGM can be formalized as:
@@
$\mathcal{Z}_{ref}=\mathrm{Softmax}($ $\frac{\mathrm{TAR}(W_{q}\mathcal{Z}_{t},W_{k}\mathcal{T}_{z})}{\sqrt{d_{k}}}W_{v}(\mathcal{T}_{z}),$ (4)
where $W_{q}$, $W_{k}$, and $W_{v}$ represent the linear layers for query, key, and value in original cross-attention, respectively. The hyper-parameter $d_{k}$ is acquired from the query dimensions. $\mathrm{TII}(\cdot,\cdot)$ and $\mathrm{TAR}(\cdot)$ denotes the functions of TII and TAR. In the end, we get refined noisy latent feature ${Z}_{ref}$. A detailed description of these three modules is provided as follows.
@@
In previous work Guo et al. (2023b); Girdhar et al. (2023), the text embedding $\mathcal{T}_{rep}$ is repeated equally $f$ times, resulting in $\mathcal{T}_{rep}\in\mathbb{R}^{f\times n\times c}$, $n$ denoting the length of the embedding vector. We inject temporal information into the embedding before performing spatial cross-attention, thereby enabling distinct focal points on the text within different frames. In Temporal Information Injector (TII), we initially reshape the noisy latent $\mathcal{Z}$ from $\mathbb{R}^{f\times h\times w\times c}$ to $\mathbb{R}^{(hw)\times f\times c}$ and apply temporal self-attention to acquire $\mathcal{Z}_{t}$. Then, we conduct spatial cross-attention, using the repeated text embedding $\mathcal{T}_{rep}$ as queries and the noisy latent $\mathcal{Z}_{t}\in\mathbb{R}^{f\times(hw)\times c}$ as both keys and values, resulting in the text embedding $\mathcal{T}_{z}$ with frame-specific temporal information. The formalization of the TII module can be expressed as follows:
@@
Figure 4: Qualititive analysis. We compare the video generation results from AnimateDiff Guo et al. (2023b), DynamiCrafter Xing et al. (2023), Pika PikaLabs (2024), Gen-2 Runway (2024), and our FancyVideo.
4 Experiments
In the quantitative experiments, FancyVideo utilizes the T2I base model to generate images as the first frame. In the qualitative experiments, for aesthetic purposes and to remove watermarks, an external model is used to generate a beautiful first frame.
4.1 Qualitative Evaluation
We choose AnimateDiff Guo et al. (2023b), DynamiCrafter Xing et al. (2023), and two commercialized products, Pika PikaLabs (2024) and Gen2 Runway (2024), for a composite qualitative analysis. It is worth noting that in the quantitative experiments, the first frame of FancyVideo is generated by SDXL to achieve a more aesthetically pleasing result and to minimize the appearance of watermark (although subsequent frames may still exhibit it).
As shown in Fig. 4, our approach exhibits superior performance, outperforming previous methods regarding temporal consistency and motion richness. In contrast, AnimateDiff, DynamiCrafter, and Gen2 generate videos with less motion. Pika struggles to produce object-consistent and high-quality video frames. Remarkably, our method can accurately understand the motion instructions in the text prompt (e.g., ”A teddy bear walking … beautiful sunset.” and ”A teddy bear running … City.” case).
4.2 Quantitative Evaluation
For a comprehensive comparison with the SOTA methods, we adopt three popular benchmarks (e.g., EvalCrafter Liu et al. (2023), UCF-101 Soomro et al. (2012), and MSR-VTT Xu et al. (2016) ) and human evaluation to evaluate the quality of video generation. Among them, EvalCrafter is a relatively comprehensive benchmark for video generation currently. UCF-101 and MSR-VTT are benchmarks commonly used in previous methods Girdhar et al. (2023); Zhang et al. (2023). Meanwhile, human evaluation can compensate for the inaccuracies in existing text-conditioned video generation evaluation systems.
Dimensions Metrics Pika Gen2 Show-1 Lumiere DynamiCrafter AnimateDiff FancyVideo
VQAA($\uparrow$) 59.09 59.44 23.19 40.06 74.56 65.94 85.78
Video VQAT($\uparrow$) 64.96 76.51 44.24 32.93 59.48 52.02 74.56
Quality IS($\uparrow$) 14.81 14.53 17.65 17.64 18.37 16.54 17.38